# Vision Transformers on the Edge: A Comprehensive Survey of Model Compression and Acceleration Strategies

### v1

@@ 2.1.1 Types
Table 2: Results of different pruning techniques proposed for Vision transformers. $\downarrow$’ denotes reduction from the baseline and $\uparrow$’ denotes increase rate from the baseline models.
Algorithm Method Models Baseline Results
GFlops Params(M) Top-1 (%)
Channels
pruning  [35]
Learn dimension-wise
important score
VTP DeiT-B
@@
($\downarrow$ 45.3%)
$\downarrow$ 47.3
92.58
($\downarrow$ 1.92%)
VTP DeiT-B
10.0
($\downarrow$ 43.1%)
$\downarrow$ 48.0 $\downarrow$ 1.1%
Width & Depth Pruning  [36] Set of learnable pruning -related parameters for width pruning & shallow classifiers using intermediate information of the transformer blocks WDPruning DeiT-T
@@
13.3
($\downarrow$ 37.6%)
70.34
($\downarrow$ 1.86%)
DeiT-S
@@
3.5
($\downarrow$ 35.2%)
78.38
($\downarrow$ 1.42%)
DeiT-B
9.90
($\downarrow$ 43.4%)
55.3
($\downarrow$ 35.0%)
80.76
($\downarrow$ 1.04%)
Swin-S
6.3
($\downarrow$ 27.6%)
32.8
($\downarrow$ 30.6%)
81.80
($\downarrow$ 1.20%)
Multi- dimensional pruning  [37] Dependency based pruning criterion & an efficient Gaussian process search Multi- dimensional DeiT-S
2.9
($\downarrow$37%)
79.9
($\downarrow$0.1%)
DeiT-B
11.2
($\downarrow$36%)
82.3
($\downarrow$0.5%)
T2T-
ViT-14
2.9
($\downarrow$40%)
81.7
($\downarrow$0.2%)
Pruning the network model  [38] Single-path ViT pruning based on the token score SPViT Swin-S
6.35
($\downarrow$ 26.4%)
82.71
($\downarrow$ 0.49%)
Swin-T
3.47
($\downarrow$ 23.0%)
80.70
($\downarrow$ 0.50%)
PiT-S
2.22
($\downarrow$ 23.3%)
80.38
($\downarrow$ 0.58%)
PiT-XS
1.13
($\downarrow$ 18.7%)
77.86
($\downarrow$ 0.24%)
Patch pruning  [39] Layer-by-layer top down pruning PS-ViT DeiT-T
0.7
($\downarrow$ 46.2%)
72.0
($\downarrow$ 0.20%)
DeiT-S
@@
($\downarrow$ 40.4%)
81.1
($\downarrow$ 0.40%)
Structural pruning  [40] Prune the head number & head dimensions inside each layer UVC DeiT-T
0.51
(39.12%)
70.6
($\downarrow$ 1.6%)
DeiT-S
2.32
(50.41%)
78.82
($\downarrow$ 0.98%)
DeiT-B
@@
(45.50%)
80.57
($\downarrow$ 1.23%)
T2T-
ViT-14
2.11
($\downarrow$ 44.0%)
78.9
($\downarrow$ 2.6%)
Global structural pruning  [41] Latency-aware, Hessian- based importance- based criteria
NViT-B
$+$ ASP
DeiT-B
@@
(2.57×2.57$\times$2.57 ×)
17
(5.14$\times$)
83.29
($\downarrow$ 0.07%)
NViT-H
$+$ ASP
Swin-S
@@
(2.85×2.85$\times$2.85 ×)
15
(5.68$\times$)
82.95
($\downarrow$ 0.05%)
NViT-S
$+$ ASP
DeiT-S
4.2
(4.24$\times$)
10.5
(8.36$\times$)
82.19
($\uparrow$ 1.0%)
NViT-T
$+$ ASP
DeiT-T
1.3
(13.55$\times$)
3.5
(24.94$\times$)
76.21
($\downarrow$ 1.71%)
Collaborative pruning  [42] Structural pruning on MSA attention & FFN by removing unnecessary parameter groups SAViT DeiT-B
@@
($\downarrow$40.1%)
82.75
($\uparrow$ 0.91%)
DeiT-S
3.1
($\downarrow$31.7%)
14.7
($\downarrow$33.5%)
80.11
($\uparrow$ 0.26%)
DeiT-T
0.9
($\downarrow$24.4%)
@@
($\downarrow$25.2%)
70.72
($\downarrow$ 1.48%)
Structured sparse pruning  [43] Removing sub-modules like self-attention heads by manipulating weight, activation,& gradient S2ViTE-B DeiT-B
11.8
($\downarrow$33.13%)
@@
($\downarrow$34.4%)
82.22
($\uparrow$ 0.38%)
S2ViTE-S DeiT-S
3.1
($\downarrow$31.7%)
14.6
($\downarrow$31.63%)
79.22
($\downarrow$ 0.63%)
S2ViTE-T DeiT-T
0.9
($\downarrow$23.69%)
@@
($\downarrow$26.3%)
70.12
($\downarrow$ 2.08%)
Bottom-up cascade pruning [44] Token pruning & channel pruning using a hyperparameter from one to last block VTC-LFC DeiT-B $\downarrow$54.4%
@@
($\downarrow$ 34.25%)
81.6
($\downarrow$ 0.20%)
DeiT-S $\downarrow$47.1%
15.3
($\downarrow$ 30.77%)
79.6
($\downarrow$ 0.20%)
DeiT-T $\downarrow$41.7%
4.2
($\downarrow$ 26.32%)
71.0
($\downarrow$ 1.20%)
Cascade ViT pruning [45] Utilizing the sparsity to prune PH-regions in MSA & FFN progressively & dynamically CP-ViT ViT-B $\downarrow$46.34-
76.75
($\downarrow$ 1.16%)
DeiT-B $\downarrow$41.62-
81.13
($\downarrow$ 0.69%)
CP-ViT ViT-B $\downarrow$29.03-
96.20
($\downarrow$ 1.93%)
DeiT-B $\downarrow$30.08-
98.01
($\downarrow$ 1.09%)
CP-ViT ViT-B $\downarrow$32.05-
84.79
($\downarrow$ 2.34%)
DeiT-B $\downarrow$30.92-
89.68
($\downarrow$ 1.17%)
Figure 3: Static Pruning vs Dynamic Pruning
In contrast, dynamic pruning applies during the runtime based on specific criteria, such as the importance of neurons or weights. One of the significant drawbacks of static pruning is that it relies on a fixed pruning schedule and rate, which is determined before training begins. That means that static pruning does not adjust to the network’s learning progress or the changing importance of neurons or weights during training. Dynamic pruning evaluates and adjusts the pruning criteria during training based on real-time importance assessments to overcome the limitation of static pruning. Most of the recent pruning techniques use dynamic pruning techniques [45, 13, 38] to get the accuracy without loss of any information.
2.1.2 Pruning Techniques for Vision Transformer
Pruning methods for ViT-based models remain an underexplored area, with only a handful of studies in recent years. This section provides a brief overview of the current SOTA pruning techniques for ViTs.
Important-based Pruning Zhu et al. [35] pioneered a ViT pruning approach that removes dimensions with lower importance scores, achieving a high pruning ratio without sacrificing accuracy. Their study observed that a significant portion of ViT’s computational cost comes from multi-head self-attention (MSA) and multi-layer perceptron (MLP). To address this, they introduced visual transformer pruning (VTP)—the first dedicated pruning algorithm for ViTs. VTP operates in three key steps: (1) L1 sparse regularization is applied during training to identify less significant channels, (2) channel pruning eliminates redundant computations, and (3) finetuning. This VTP approach managed to preserve the robust representative capability of the transformer while reducing the model’s computational cost. Another recent study [36] utilized learning a unique saliency score and threshold for each layer to implement width pruning. This learning saliency score allows for a more effective, non-uniform allocation of sparsity levels across different layers. Additionally, The proposed model utilized supplementary plug-in classifiers to prune the transformer’s trailing blocks. This approach enabled the construction of a sequential variant of the pruned model, capable of removing blocks within a single training epoch, thereby simplifying the control of the trade-off between the network’s performance and the rate of pruning  [36]. Moreover, another study by Tang et al. [39] proposed a patch-slimming approach that reduced unimportant patches in a top-down manner. The authors calculated their importance scores for the final classification feature to identify unimportant patches. The proposed method also identified the important patches in the last layer of the blocks and then utilized them to select the previous layer patches.
Token Pruning Kong et al. [38] introduced a latency-aware soft token pruning framework, SP-ViT. This framework was implemented on vanilla transformers such as data-efficient image transformers (DeiT) [47] and swin transformers [6]. The authors proposed a dynamic attention-based multi-head token selector for adaptive instance-wise token selection. Later, they incorporated a soft pruning method that consolidated less informative tokens into a package token instead of entirely discarding them identified by the selector module. The authors deployed their proposed method on ImageNet-1k with baseline models Swin-S, Swin-T, PiT-S, and PiT-Xs.
Structure Pruning Recently, many studies on ViT pruning have embraced structure pruning techniques to optimize model efficiency. Yu et al. [40] proposed a structure pruning in a ViT named UVC where they pruned the head’s number and dimension inside each layer. Experiments of this paper were conducted in various ViT models (e.g., DeiT-Tiny and T2T-ViT-14) on ImageNet-1k [5] datasets. DeiT-Tiny [47] cut down to 50% of the original FLOPs while not dropping accuracy much in this study. Another study [42] proposed structure pruning on MSA attention and feedforward neural network (FFN) by removing unnecessary parameter groups. Other studies on structure pruning named NViT [41] proposed hessian-based structure pruning criteria comparable across all layers and structures. Moreover, it incorporated latency-aware regularization techniques to reduce latency directly. Another study on the structure pruning in ViT called S2ViT [43] removed submodules like self-attention heads by manipulating the weight, activations & gradients.
Cascade Pruning Cascade pruning combines multiple pruning techniques to reduce parameters and GFLOPs while preserving accuracy. A standout method named VTC-LFC [44] aimed to improve the identification of informative channels and tokens in a model, leading to better accuracy preservation. This approach introduced a bottom-up cascade (BCP) pruning strategy that gradually prunes tokens and channels, starting from the first block and advancing to the last. The pruning process is controlled by a hyper-parameter called a global allowable drop, ensuring the performance drop remains within an acceptable range. Additionally, BCP guarantees efficient compression without sacrificing model performance by pruning each block and immediately stopping the compression process when the performance drop reaches a predefined threshold. Another cascade ViT pruning  [45] utilized the sparsity for pruning PH-regions (containing patches and heads) in the MSA & FFN progressively and dynamically. The authors conducted experiments on three different types of datasets: ImageNet-1k [5], CIFAR-10 [48], and CIFAR-100 [48].
@@ 2.1.3 Discussion
Pruning is utilized as a fundamental way to reduce the computation of the pre-trained ViT models. For ViT, the development of the pruning methods has systematically covered each perspective of model design, making the current pruning methods more flexible and well-organized for ViT models. We summarized all the core information about pruning techniques on ViT in comparison to GFlops reductions with the percentage of reduction from baseline, parameters reductions, and top-1 accuracy in Table 2. Top-1 accuracy shows the accuracy loss with the proposed methodology from baseline backbone ViT architecture. Recent pruning studies on ViT models, as shown in Table 2, have predominantly focused on the ImageNet-1k dataset [51], with the exception of CP-ViT [45], which conducted experiments on the CIFAR dataset [48]. However, the training/finetuning cost is one of the critical points to consider in the hardware-inefficient ViT models in the pruning methods. Therefore, training-efficient or fine-tuned free pruning techniques need more attention in the near future for efficient deployment on the edge. This necessitates a more precise estimation of parameter or block sensitivity using limited data, as well as a deeper exploration of the information embedded within hidden features during training and inference.
Table 3: Results of different KD (classification) techniques proposed for vision transformers. Here, MSE loss means mean square error loss.
@@
Algorithm Method Loss function Dataset Teachers Students
@@
Fine-grained manifold  [52] Patch-level manifold space method Manifold distillation loss (MD Loss) ImageNet-1k [51] CaiT-S24 83.4% DeiT-T
@@
Target-aware Transformer  [17] One-to-all spatial matching KD Vanilla distillation + L2 + Task loss ImageNet-1k [51] ResNet34 72.4% ResNet18 72.1 ($\uparrow$+2.0%)
Cross Inductive Bias Distillation  [53] Co-advising the student models with lightweight teacher model Kull back divergence + Cross entropy loss ImageNet-1k [51] ResNet18 83.4% Transformer-Ti 88.0 ($\uparrow$+1.5%)
Attention probe  [54] Probe distillation & Knowledge distillation Probe distillation + cross-entropy CIFAR-100 DeiT-XS 76.30% DeiT-XTiny
71.82
($\uparrow$+6.36%)
CIFAR-10 DeiT-XS 96.65% DeiT-XTiny
93.95
($\uparrow$+7.64%)
MNIST DeiT-XS 99.39% DeiT-XTiny
99.07
($\uparrow$+0.01%)
MiniViT  [55] Weight distillation Self-attention distillation + Hidden-state distillation + prediction loss ImageNet-1k [51] RegNet- 16GF 82.9% DeiT-B 83.2 ($\uparrow$+1.4%)
TinyViT  [56] Reusing the teachers’ prediction & data augmentation for student Cross entropy loss ImageNet-1k [51] CLIP- ViT-L 84.8% Swin-T
@@
($\uparrow$+2.2%)
DeiT-S 82.0 ($\uparrow$+2.1%)
@@
($\downarrow$+1.0%)
2.2 Knowledge Distillation
Knowledge distillation (KD) is another model compression technique in machine learning where a smaller model (the "student") is trained to reproduce the behavior of a larger model (the "teacher"). The purpose is to transfer the "knowledge" from the larger model to the smaller one, thereby reducing computational resources without significantly losing accuracy. Pruning is a direct way to reduce the complexity of the original model, whereas KD involves training a new, more compact model that is easy to deploy. By using KD during finetuning, the pruned model can benefit from the insights and information captured by the larger teacher model. KD helps to regain even surpass the original performance and compensate for the accuracy loss during the other compression techniques (e.g., pruning, quantization).
2.2.1 KD techniques for vision transformer
Touvron et al. [47] leveraged KD to train the transformer with a significantly smaller dataset than the traditionally required dataset. The authors introduced a distillation token, an additional learnable vector used alongside the class token during training. The proposed method achieved 84.5% top-1 accuracy on the ImageNet-1k [4] dataset, requiring fewer training data resources and computing power than ViT without KD.
Another study by Hao et al.  [52] utilized every patch information to introduce a fine-grained manifold distillation method. In the manifold distillation method, the authors considered ViT as a feature projector that sets image patches into a sequence of manifold space layer by layer. The authors then teach the student layers to generate output features having the same patch-level manifold structure as the teacher layer for manually selected teacher-student layers. These output features are normalized and reshaped to compute a manifold relation map, a representation of the manifold structure of the features. However, the manifold relation map computation is resource-consuming and needs to simplify the computation. To solve the computational issue, the authors decoupled the manifold relation map into an intra-image relation map, an inter-image relation map, and a randomly sampled relation map. The authors utilized manifold distillation loss (MD Loss), the sum of individual loss from all three decoupled manifold relation maps.
Table 4: Results of different KD (object detection) techniques proposed for vision transformers.
@@ 2.2.1 KD techniques for vision transformer
Fine-grained manifold  [52] ImageNet-1k [51] COCO-2017 Swin-T 43.7 Swin-T 44.7($\uparrow$1.0)
@@
Moreover, Lin et al.  [17] proposed a one-to-all spatial matching KD technique. The proposed approach involves distilling the knowledge from each pixel of the teacher feature to all spatial locations of the student features based on their similarity. This similarity is determined using a target-aware transformer. By leveraging this target-aware transformer, the teacher’s feature information can be effectively transferred and distilled to different spatial locations in the student’s features.
Furthermore, Wang et al.  [54] proposed another KD method called attention probes. The main idea of this paper was to streamline ViTs through a two-step process using unlabelled data gathered from varied sources. The authors developed an ’attention probe’ in the initial phase to discern and select significant data. The authors then used the selected data to instruct a compact student transformer by applying a probe-based KD algorithm. This algorithm was designed to optimize the resemblance between the resource-intensive teacher model and the more efficient student model, considering both the final outputs and intermediate features. The proposed method used cross-entropy (CE) and probe distillation functions for distilling intermediate features for calculating loss. Another study named DearKD  [57] proposed the KD methods on self-generative data and used representational KD on intermediate features with response-based KD. The proposed paper used mean square error (MSE) distillation loss for hidden features, CE loss for hard label distillation, and intra-divergence distillation loss function to calculate the loss. It was noteworthy to see DearKD surpass the performance of the baseline transformer that trained with all ImageNet data, even though it only used 50% of the data. The authors then evaluated the proposed technique on the ImageNet dataset and achieved 74.8% top-1 accuracy in a tiny version, which is 2% better than the DeiT-Tiny  [47]. Moreover, TinyViT [56] highlights that smaller ViTs can benefit from larger teacher models trained on extensive datasets, such as distilling the student model on ImageNet-21k and finetuning on ImageNet-1k. To optimize computational memory, TinyViT introduced a strategy that pre-stores data augmentation details and logits for large teacher models, reducing memory overhead. Additionally, MiniViT [55] argues by introducing a weight multiplexing strategy to reduce parameters across consecutive transformer blocks. Moreover, they employed weight distillation on self-attention mechanisms to transfer knowledge from large-scale ViT models to the smaller, weight-multiplexed MiniViT models.
Table 5: Results of different KD (semantic segmentation) techniques proposed for vision transformers.

### v2

@@ 2.1.2 Pruning Techniques for Vision Transformer
Table 4: GFLOPS reduction comparison of pruning techniques across different ViT models. $\downarrow$’ denotes the reduction rate from the baseline ViT model.
Algorithm GFLOPS ($\downarrow$)
DeiT-T DeiT-S DeiT-B Swin-S T2T-ViT-14
Baseline 1.3 4.6 17.6 8.7 4.8
VTP [38] - -
@@
($\downarrow$ 43.2%)
WDPruning [40]
0.7
($\downarrow$ 46.2%)
@@
9.9
($\downarrow$ 43.8%)
6.3
($\downarrow$ 27.6%)
MD [51] -
2.9
($\downarrow$ 37.0%)
11.2
($\downarrow$ 36.4%)
2.9
($\downarrow$ 39.6%)
SPViT [37] - - -
6.4
($\downarrow$ 26.4%)
PS-ViT [39]
@@
($\downarrow$ 35.4%)
UVC [41]
0.5
($\downarrow$ 61.5%)
2.3
($\downarrow$ 50.0%)
@@
($\downarrow$ 54.5%)
2.1
($\downarrow$ 56.0%)
NViT [42]
1.3
($\downarrow$ 0.0%)
4.2
($\downarrow$ 8.7%)
@@
($\downarrow$ 61.4%)
@@
($\downarrow$ 28.7%)
SAViT [47]
0.9
($\downarrow$ 30.8%)
3.1
($\downarrow$ 32.6%)
@@
S2ViTE [43]
0.9
($\downarrow$ 30.8%)
3.1
($\downarrow$ 32.6%)
11.8
($\downarrow$ 33.0%)
VTC-LFC [44]
0.76
($\downarrow$ 41.7%)
2.4
($\downarrow$ 47.1%)
8.03
($\downarrow$ 54.4%)
CP-ViT [36]
0.74
($\downarrow$ 43.3%)
2.7
($\downarrow$ 42.2%)
10.28
($\downarrow$ 41.6%)
Table 5: Parameter (Million) reduction comparison of pruning techniques across different ViT models. $\downarrow$’ denotes parameter reduction rate from the baseline ViT model.
Algorithm Params (M $\downarrow$)
DeiT-T DeiT-S DeiT-B Swin-S
Baseline 5.6 22 86 50
VTP
 [38]
47.3
($\downarrow$ 45.0%)
WDPruning
 [40]
3.5
($\downarrow$ 37.5%)
13.3
($\downarrow$ 39.5%)
55.3
($\downarrow$ 35.7%)
32.8
($\downarrow$ 34.4%)
NViT [42]
3.5
($\downarrow$ 37.5%)
10.5
($\downarrow$ 52.3%)
17.0
($\downarrow$ 80.3%)
15.0
($\downarrow$ 70.0%)
SAViT [47]
4.2
($\downarrow$ 25.0%)
14.7
($\downarrow$ 33.2%)
@@
($\downarrow$ 39.6%)
S2ViTE [43]
@@
($\downarrow$ 25.0%)
14.6
($\downarrow$ 33.6%)
@@
($\downarrow$ 34.0%)
VTC-LFC [44]
@@
($\downarrow$ 25.0%)
15.3
($\downarrow$ 30.4%)
@@
($\downarrow$ 34.0%)
@@ 2.1.3 Discussion
Pruning is utilized as a fundamental way to reduce the computation of the pre-trained ViT models. For ViT, the development of the pruning methods has systematically covered each perspective of model design, making the current pruning methods more flexible and well-organized for ViT models. We provide a comparative summary of these pruning methods in terms of Top-1 accuracy (Table 3), GFLOPS reduction (Table 4), and parameter reduction in millions (Table 5), comparable to baseline models. We inherit the baseline results from the official implementations of the respective models, as detailed in the tables, and compute the corresponding accuracy drop, parameter reduction, and GFLOPs reduction rates to evaluate the impact of pruning. Most pruning methods evaluated in Tables 35 focus on the ImageNet-1K dataset [45], with CP-ViT [36] being a notable exception, tested on CIFAR [52] dataset. As shown in Table 3, NViT [42] consistently achieves competitive top-1 (%) accuracy across DeiT-Tiny (T), Small (S), Base (B), and Swin-S, with a minimal drop of only 5.55% from the baseline. Interestingly, few studies(e.g, SPViT, VTP) have explored pruning across diverse ViT models such as Swin or T2T-ViT-14, highlighting a gap in the generalizability of current techniques. In terms of computational savings, UVC [41] achieves the highest GFLOPs reduction for DeiT-T/S variants. Notably, NViT also delivers substantial reductions in GFLOPs for more complex models, such as DeiT-B, while improving Top-1 accuracy—a noteworthy exception to the typical trade-off where aggressive pruning often leads to performance loss. Additionally, both WDPruning [40] and NViT [42] achieved comparable parameter sizes (see Table 5) when applied to DeiT-T. However, NViT demonstrated a more substantial reduction in the number of parameters than other ViT models. While these results demonstrate notable progress in pruning ViTs, the training and finetuning costs remain a major limitation, particularly for deployment on resource-constrained edge devices. Therefore, training-efficient or finetuned free pruning techniques need more attention in the near future for efficient deployment on the edge. This necessitates a more precise estimation of parameter or block sensitivity using limited data, as well as a deeper exploration of the information embedded within hidden features during training and inference.
(a) A schematic illustration of the logit-based KD technique.
(b) A schematic illustration of the relation-based KD technique.
(c) A schematic illustration of the feature-based KD technique.
Figure 6: An illustration depicting the three types of KD: logit-based, relation-based, and feature-based between teacher and student models.
2.2 Knowledge Distillation
Knowledge distillation (KD) is another model compression technique in machine learning where a smaller model (the "student") is trained to reproduce the behavior of a larger model (the "teacher"). The purpose is to transfer the "knowledge" from the larger model to the smaller one, thereby reducing computational resources without significantly losing accuracy. Pruning is a direct way to reduce the complexity of the original model, whereas KD involves training a new, more compact model that is easy to deploy. By using KD during finetuning, the pruned model can benefit from the insights and information captured by the larger teacher model. KD helps to regain the original performance and compensate for the accuracy loss during the other compression techniques (e.g., pruning, quantization).
2.2.1 KD techniques for Vision Transformer
Current KD techniques on ViT models can be broadly classified into three types: logit-based [55, 56], relation-based [57, 17], and feature-based [58, 59, 60] methods. Feature-based KD enables the student model to capture richer semantic representations by aligning intermediate features across hidden layers. Both logit-based and feature-based KD typically focus on transferring knowledge from individual data samples. In contrast, relation-based KD extends this idea by modeling relational knowledge across the whole dataset, aiming to capture inter-sample relationships. Figure 6 illustrates a schematic overview of these KD types, providing a visual summary of their key differences. However, A detailed analysis of these three categories has also been conducted by Yang et al.[61]. In the following subsections, we provide a detailed discussion of the current proposed techniques on ViT models of each category.
Table 6: Results of KD techniques applied to ViT models for classification tasks. MSE refers to Mean Squared Error loss. $\uparrow$’ in the student model indicates the performance improvement over the baseline student model.
@@ 2.2.1 KD techniques for Vision Transformer
Type Algorithm Method Loss function Dataset Teachers Students
@@
Relation TAT [17] One-to-all spatial matching KD Vanilla distillation + L2 loss ImageNet-1k [45] ResNet34 72.4% ResNet18 72.1 ($\uparrow$ 2.0%)
Co-advise [57] Co-advised by lightweight teachers Kull back divergence + Cross entropy loss ImageNet-1k [45] ResNet18 83.4% CiT-Ti 88.0 ($\uparrow$ 1.5%)
Feature Manifold  [58] Patch-level manifold space method Manifold distillation loss ImageNet-1k [45] CaiT-S24 83.4% DeiT-T
@@
AProbe  [60] Probe & Knowledge distillation Probe distillation + Cross-entropy loss
CIFAR-100 [52]
DeiT-XS 76.30% DeiT-XT
71.8
($\uparrow$ 6.36%)
CIFAR-10 [52]
96.65%
93.9
($\uparrow$ 7.64%)
MNIST [62]
99.39%
99.1
($\uparrow$ 0.01%)
MiniViT  [59] Weight distillation Self-attention + Hidden-state + Prediction loss ImageNet-1k [45] RegNet- 16GF 82.9% DeiT-B 83.2 ($\uparrow$ 1.4%)
Logit TinyViT  [55] Reusing teacher predictions & data augmentation Cross entropy loss ImageNet-1k [45] CLIP- ViT-L 84.8% Swin-T
@@
($\uparrow$ 2.2%)
DeiT-S 82.0 ($\uparrow$ 2.1%)
@@
($\downarrow$ 1.0%)
  • Co-advised [57]: Transformer-Ti is denoted as CiT-Ti in our table.
Logit-based KD Logit-based KD (see Figure 6(a)) focuses on learning knowledge from the last layer and aims to align the final predictions between the teacher and student. Touvron et al. [46] pioneered a benchmark KD technique for ViTs by enabling training with substantially smaller datasets. The authors introduced a distillation token, an additional learnable vector used alongside the class token during training. Recently, TinyViT [55] highlighted that smaller ViTs could benefit from larger teacher models trained on extensive datasets, such as distilling the student model on ImageNet-21k and finetuning on ImageNet-1k. This approach reduced the computation significantly during the training of the student model. To optimize computational memory, TinyViT introduced a strategy that pre-stores data augmentation details and logits for large teacher models, reducing memory overhead. One problem with TinyViT is that teacher models still use large datasets. To solve this issue, DearKD  [56] proposed the KD methods on self-generative data and used representational KD on intermediate features with response-based KD. The proposed paper used mean square error (MSE) distillation loss for hidden features, CE loss for hard label distillation, and intra-divergence distillation loss function to calculate the loss. It was noteworthy that DearKD surpassed the performance of the baseline ViT model trained with ImageNet-21K datasets, even though it only used 50% of the data. However, logit-based KD only transfers output-level predictions, neglecting intermediate representations such as features or attention. This shallow supervision limits its effectiveness, especially when the student and teacher architectures differ significantly.
Table 7: Results of KD techniques applied to ViT models for object detection tasks.
@@
Manifold  [58] ImageNet-1k [45] COCO-2017 Swin-T 43.7 Swin-T 44.7($\uparrow$1.0)
@@
Feature-based KD As we mentioned above, logit-based KD neglects intermediate-level supervision for complete guidance. To solve this problem, feature-based KD (see Figure 6(c)) focuses on exploring intermediate feature information, such as feature maps and their refined information. Hao et al. [58] introduced a fine-grained manifold distillation technique, which considered ViT as feature projectors that map image patches into a sequence of high-dimensional manifold spaces. The authors then teach the student layers to generate output features having the same patch-level manifold structure as the teacher layer for manually selected teacher-student layers. These output features are normalized and reshaped to compute a manifold relation map, a representation of the manifold structure of the features. However, the manifold relation map computation is resource-consuming and needs to be simplified. To address this computational issue, the authors proposed a decomposition strategy that separates the manifold relation map into three components: intra-image, inter-image, and randomly sampled relation maps. A dedicated manifold distillation loss (MD Loss) is then applied by aggregating the losses computed from each decoupled manifold relation map. However, the proposed method only relies on prediction-level distillation. To consider the other level (e.g., attention level), Minivit [59] considered both attention level and hidden-state distillation, guiding a small model to replicate the behavior of the teacher model. The key idea of the proposed method is to multiplex the weights of consecutive transformer blocks by sharing the weights of the MSA and MLP layers. Besides, weight distillation over self-attention transfers knowledge from a large teacher model to a compact student. To ensure effective knowledge transfer, MiniViT introduced a multi-level distillation loss that combines three components: (i) a prediction-logit loss that aligns the student’s output with the teacher’s softened logits, (ii) an attention distillation loss that encourages similarity in self-attention, and (iii) a hidden-state distillation loss that preserves structural relationships. Furthermore, Wang et al. [60] introduced another feature-based KD technique for ViT named attention probes (AProbes), which streamline ViTs using unlabeled data. The proposed method operates in two stages: first, an attention-based probing mechanism identifies informative samples from diverse unlabeled datasets. Next, a probe-guided distillation process transfers knowledge to the student model by aligning the final predictions and intermediate features with the teacher. Both stages are achieved through cross-entropy loss and a dedicated probe distillation loss. However, feature-based KD methods only focus on individual sample representations despite leveraging intermediate features and often overlook the underlying dependencies across layers. This limitation restricts their ability to capture the holistic knowledge embedded in the teacher model fully. To overcome this limitation, relation-based KD aims to model and transfer relational information.
Relation-based KD Relation-based KD (see Figure 6(b)) is often modeled as an instance graph, where nodes represent sample embeddings and edges are weighted by similarity metrics. One notable example is Co-advise [57], which introduced a multi-teacher framework where lightweight teacher models with diverse inductive biases (e.g., CNNs) collaboratively guide a student ViT model. The key insight is that the inductive bias of a teacher has a more significant impact on student performance than their standalone accuracy. By distilling complementary inductive biases, the student gains a more comprehensive representation of the data. A token-level alignment mechanism is also proposed to match the inductive bias of student tokens with their respective teachers. Another method, target-aware transformer (TAT) [17], addressed spatial misalignment in distillation. It transferred each pixel in the teacher feature map to all spatial locations in the student using a transformer-based similarity mechanism, enabling fine-grained relational alignment. While Co-advise emphasizes architectural diversity and inductive bias, and TAT focuses on spatial adaptability, both exemplify the strength of relation-based KD in enhancing knowledge transfer beyond isolated feature matching.
Table 8: Results of KD techniques applied to ViT models for segmentation tasks.